Suppose that Lil John Industries’ equity is currently selling for $31 per share and that 1.4 million shares are outstanding. Assume the firm also has 24,000 bonds outstanding, and they are selling at 102 percent of par.

What are the firm’s current capital structure weights?
Answer #1

Capital structure weights are computed based on market value.

Market value of equity = Shares are outstanding * Selling price per share

= 1.4 million * 31

= 43,400,000

Market value of Debt =  bonds outstanding * Selling price per bond

= 24000*1000*102%

= 24,480,000

Weight of equity = Market value of equity / (Market value of equity+Market value of Debt)

= 43400000/(43400000+24480000)

= 43400000/67880000

= 0.63936358279

= 0.6394

= 63.94%

Weight of debt = Market value of debt / (Market value of equity+Market value of Debt)

= 1 - Weight of equity

= 1-.6394

= 0.3606

= 36.06%

note: It is general practice to take $1,000 as face value of bonds when no details are given.